Ok heres my question . ===> Insert parentheses in the expression 6 + 10 x 2 so that the expression equals ( 12 + 1 ) x 2 .
THANX SOO MUCH! I REALLY NEED HELP THANKS SO MUCH

Answer by jim_thompson5910(35256)

( 12 + 1 ) x 2 = ( 13 ) x 2

( 12 + 1 ) x 2 = 26

Since 6 + 10 x 2 = 26 already, no parenthesis are needed, but if you want, you can say

6 + (10 x 2) = 26